# UniMo: Universal Motion Correction For Medical Images without Network Retraining

### v2

@@ Abstract
In this paper, we introduce a Universal Motion Correction (UniMo) framework, leveraging deep neural networks to tackle the challenges of motion correction across diverse imaging modalities. Our approach employs advanced neural network architectures with equivariant filters, overcoming the limitations of current models that require iterative inference or retraining for new image modalities. UniMo enables one-time training on a single modality while maintaining high stability and adaptability for inference across multiple unseen image modalities. We developed a joint learning framework that integrates multimodal knowledge from both shape and images that faithfully improve motion correction accuracy despite image appearance variations. UniMo features a geometric deformation augmenter that enhances the robustness of global motion correction by addressing any local deformations whether they are caused by object deformations or geometric distortions, and also generates augmented data to improve the training process. Our experimental results, conducted on various datasets with four different image modalities, demonstrate that UniMo surpasses existing motion correction methods in terms of accuracy. By offering a comprehensive solution to motion correction, UniMo marks a significant advancement in medical imaging, especially in challenging applications with wide ranges of motion, such as fetal imaging. The code for this work is available online  [^1] https://github.com/IntelligentImaging/UNIMO.
Index Terms:
Deep Learning, Medical Imaging Analysis, Motion Compensation Analysis, Multi-modality Fusion, Rigid Transformation Estimation
@@ I Introduction
Motion remains to be a challenge in medical imaging. For reliable imaging and image analysis, motion should be avoided, corrected, or compensated for, to minimize motion artifacts [1, 2, 3, 4]. Although classical image registration methods for motion correction can be effective, they often necessitate full execution of an optimization process on non-convex cost functions that are often only a surrogate measure of image alignment [5, 6, 7, 8, 9]. Thus, these techniques are typically limited by a narrow capture range, especially when addressing large motions. This constraint reduces their precision and applicability in challenging applications such as fetal MRI where subjects may move in wide angles [10]. To address these limitations, recent research has increasingly turned to deep learning based models for predictive motion estimation [11].
Deep learning-based motion correction techniques have seen widespread application across various imaging modalities, with significant developments in two primary areas: i) A major focus on Magnetic Resonance Imaging (MRI), including functional MRI (fMRI), T1, T2, and diffusion-weighted MRI, and quantitative MRI. These models are extensively studied due to their broad applicability and substantial impact in the field, covering a wide range of clinical scenarios. ii) Motion correction for other modalities, such as ultrasound and positron emission tomography (PET), and Computed Tomography (CT) where advancements and methodologies are also being explored, though to a lesser extent compared to MRI.
Learning-based motion correction in MRIs focuses on two key approaches: i) image-based motion correction and ii) techniques using k-space data. Image-based methods target motion correction by processing reconstructed MRI images with deep neural networks, such as Convolutional Neural Networks (CNNs) [12, 13] and other deep learning architectures [14, 15, 16, 17]. These approaches often employ convolutional encoder-decoder structures, leveraging downsampling and upsampling for feature extraction and reconstruction [18, 19]. Advanced models like U-Nets, recurrent networks, and transformers handle temporal dependencies and spatial relationships [20, 21, 17, 22, 14]. Additionally, prior-assisted methods improve correction by incorporating supplementary data such as multi-contrast or dynamic images, enhancing performance through architectures that integrate this information [23, 24]. These methods offer robust solutions by combining motion correction with other learning-based tasks, ultimately improving MRI reconstruction quality, e.g. [22].
@@
Despite significant advancements, state-of-the-art methods remain heavily constrained by the specific training dataset for each modality, limiting their effectiveness in correcting motion in unseen image modalities. While one approach [16] shows promise in handling motion for unseen modalities, it struggles with large motions in image pairs that are significantly contaminated by noise [38] and subject to contrast changes. For instance, substantial intensity and contrast variations in images pose significant challenges, as current algorithms largely rely on image intensities while overlooking shape information. This limitation significantly impairs the applicability of motion correction models in a multi-modality context.
Motivated by the need for high-accuracy motion correction for multiple modalities, we present a novel motion tracking approach that corrects motion artifacts without necessitating network retraining. Our method leverages the strengths of deep learning while circumventing the need for extensive retraining, thereby offering a robust and efficient solution for real-time motion correction. This innovative approach does not make restricting assumptions about the range or type of motion thus enhances the stability and accuracy of motion correction under varying imaging conditions. By eliminating the retraining requirement, our method significantly reduces computational overhead and facilitates seamless integration into clinical workflows, thus advancing the state-of-the-art in motion estimation and correction.
Our approach builds on the extensive body of work in motion correction and tracking, combining the robustness of traditional techniques with the flexibility and efficiency of modern deep learning models. By integrating shape information and leveraging advanced neural network architectures, we provide a comprehensive solution that overcomes the limitations of existing methods, ensuring accurate and reliable motion correction in real-time clinical settings. This advancement represents a significant step forward in the field of medical imaging, particularly in challenging domains such as fetal MRI, and paves the way to use advanced imaging to develop more effective diagnostic and therapeutic interventions. This paper is a major extension and advancement over our recent works [15, 14]. The major contributions of UniMo are:
  • UniMo is the first method for motion correction on any image modality without retraining neural networks;
  • UniMo offers multimodal motion correction with stable convergence in training, and real-time inference for highly accurate registration of a pair of images.
@@
  • UniMo paves a way for real-time image reconstruction where images are severely corrupted, for example in slice-to-volume registration using stack of slices with unpredictable inter-slice motion artifacts.
@@ II-A Rigid Motion Estimation via Equivariant Filters
Rather than estimating the transformation function $Q$ directly in the original image space, efficient approaches have demonstrated that it can be computed using low-dimensional representations, such as key points or equivariant features of images [16]. In our approach, we utilize equivariant filters due to their proven stability and accuracy in single-modality motion correction. The closed-form solution for both translation and rotation parameters is:
@@
where $\bar{S}$, $\bar{T}$ represent the low-dimensional representations of the source and target respectively. ${U}\Sigma V^{*}=\bar{S}\cdot\bar{T}^{T}$, $U$ and $V^{*}$ are real orthogonal matrices, $\Sigma$ is a diagonal matrix with non-negative real numbers on the diagonal. We set the determinant of the $\mathcal{R}$ in Eq. (2) to $1$ guarantees that it accurately reflects a rigid transformation.
@@ II-B Deformation Correction via Deformable Registration
where the distance function ${\rm Dist}(\cdot,\cdot)$ measures the dissimilarity between images. Commonly used distance functions include sum-of-squared difference of image intensities [39], normalized cross correlation [40], and mutual information [41, 42]. The regularization term (~$(\tilde{\mathcal{L}}\tilde{v}_{0},\tilde{v}_{0})$ enforces spatial smoothness of transformations with $\tilde{\mathcal{L}}$ being a symmetric and positive-definite differential operator in the Fourier domain. It converts a vector field $\tilde{v}$ to a momentum vector by $\tilde{m}=\tilde{\mathcal{L}}\tilde{v}$. The deformation $\psi_{1}$ corresponds to $\tilde{\psi}_{1}$ in Fourier space via the Fourier transform $\mathcal{F}(\psi_{1})=\tilde{\psi}_{1}$, or its inverse $\psi_{1}=\mathcal{F}^{-1}(\tilde{\psi}_{1})$.
@@
where $\star$ is a circular matrix-vector field auto-correlation [^2] Auto-correlation operates on zero-padded signals followed by truncating to the bandlimits in each dimension to ensure the output remains bandlimited.. $\tilde{\mathcal{K}}$ is a smoothing operator and $\tilde{\mathcal{D}}\tilde{v}$ is a tensor product with $\tilde{\mathcal{D}}(\mathbf{\xi})=i\sin(2\pi\mathbf{\xi})$ representing the Fourier frequencies of a central difference Jacobian matrix $D$. The operator $\tilde{\nabla}\cdot$ is the discrete divergence operator that is computed as the sum of the Fourier coefficients of the central difference operator $\tilde{D}$ along each dimension, i.e., ~$\tilde{\nabla}\cdot\mathbf{\xi}=\sum\limits_{j=1}^{d}i\sin(2\pi\mathbf{\xi}_{j})$. Since $\tilde{\mathcal{K}}$ on the left side of Eq. (4) is a low-pass filter that suppresses high frequencies in the Fourier domain, all operations are easy to implement in a truncated low-dimensional space by eliminating high frequencies.
@@ II-C Joint Optimization
$=\underbrace{\text{Dist}[S\circ Q(\mathcal{T},\mathcal{R}),T]+\text{Reg}(\mathcal{T},\mathcal{R})}_{\text{Motion Correction}}$
$+\underbrace{\text{Dist}(\mathbf{\hat{S}}\circ\psi_{1}(\tilde{v}_{0})),T)+(\tilde{\mathcal{L}}\tilde{v}_{0},\tilde{v}_{0})}_{\text{Distortion Correction}},$
@@
Such joint approaches have demonstrated enhanced accuracy and robustness for single image modalities [15, 14]. However, these frameworks fall short when image pairs exhibit significant intensity changes. The issue arises because both dissimilarity terms in Eq.(II-C) are highly dependent on the images themselves, increasing the likelihood that the optimization will be biased by intensity variations. This limitation motivated the development of a multimodal motion correction framework that integrates knowledge from both image intensities and shape to effectively reduce errors caused by image intensity variations.
@@ III Methodology
III-A Multi-modal Motion Correction
In this section, we present a multimodal motion correction approach that utilizes both image and shape information. Our objective is to develop an optimal motion correction solution, denoted as $\boldsymbol{Q}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{R}})$, which exhibits increased robustness to variations in image intensity. We first define the basic composition operation between two rigid transformations,
@@ III-A Multi-modal Motion Correction
After obtaining the multimodal rigid transformation $\boldsymbol{Q}$, for deformation estimation we estimate $\tilde{v}^{I}_{0}$ and $\tilde{v}^{G}_{0}$ (representing the transformation fields) in the spatial domain for images and segmentations respectively. The derivation of a multimodal velocity field is not the primary objective of this work. Instead, we formalize the multimodal objective function for optimization as follows,
@@
III-B Multimodal Rigid Transformation for Motion Correction
@@
    Convert $\mathbf{q}$ back to a rotation matrix $\boldsymbol{\mathcal{R}}$ (details in Appendix A);
@@
    The final multimodal rigid transformation matrix is
@@ III-C Network Design and Training
We develop a deep learning framework to estimate the objective expressed in Eq. (8). Our framework comprises two major sub-modules: i) A multimodal rigid motion correction neural network, parameterized by equivariant filters, to produce $\boldsymbol{Q}$; and ii) A multimodal deformation correction network, implemented using UNet, with deformable shape augmentation to estimate $\tilde{v}^{I}_{0}$ and $\tilde{v}^{G}_{0}$. Our framework is illustrated in Fig. 1. In the following sections, we provide a detailed description of our network architecture and the formulation of our network loss.
Figure 1: An illustration of the network architecture of our proposed motion correction learning framework, UniMO. Top: Motion correction is based on equivariant neural networks that take both images and segmentations as input. The low-dimensional spatial means of the images and segmentations are estimated simultaneously. An enhanced multimodal rigid transformation $Q$ is computed from both domains. The rigid loss function is calculated between the aligned and target images/labels. Bottom: A geometric shape augmenter is incorporated into the U-Net based neural networks. It takes both aligned and target images/segmentations and estimates the transformation fields for both domains. Unlike the motion correction module, the loss function of the geometric augmenter is separately defined for segmentations and images.
Multimodal Rigid Motion Correction Equivariant Network
Let $\Theta=(\mathcal{T}_{I},\mathcal{R}_{I},\mathcal{T}_{G},\mathcal{R}_{G})$ represent the encoder parameters that learn rigid parameters from image and shape spaces, with $Q_{I}(\Theta)$ and $Q_{G}(\Theta)$ denoting the transformation functions yielded by the learned rigid parameters from the low-dimensional representations. Let $\boldsymbol{Q}$ denote the multimodal rigid transformation. The rigid correction loss is computed between the aligned outputs and targets. Note that CNNs or recurrent neural networks for extracting low-dimensional features from images are also optional [13, 12, 20]. However, such representations are limited in capturing the true nature of rigid transformations. In our approach, we employ an efficient method to compute rigid transformations within the equivariant neural network by calculating the spatial means of images [16]. We express the general formulation for the rigid motion network, which takes source images (or segmentations) and target images (or segmentations), as follows:
@@
In addition to estimating deformations using LDDMM [45], we provide a deep learning model that learns stationary velocity fields [51] while maintaining comparable model accuracy. Advanced predictive image registration models, including TransMorph [52] and DiffuseMorph [53], can be integrated into our framework. Analogous to our previous work [14], we extend UniMo to a spatio-temporal approach that incorporates time information to handle the motion among data sequences. Please check the reference [14] for more details.
@@
The objective function for the multimodal model is defined as:

### v3

# A Unified Deep Learning Framework for Motion Correction in Medical Imaging

@@ Abstract
Deep learning has shown significant value in image registration, however, current techniques are either limited by the type and range of motion they can handle, or require iterative inference and/or retraining for new imaging data. To address these limitations, we introduce UniMo, a Unified Motion Correction framework that leverages deep neural networks to correct for diverse types of motion in medical imaging. UniMo exploits an alternating optimization scheme for a unified loss function to train an integrated model of 1) an equivariant neural network for global rigid motion correction and 2) an encoder-decoder network for local deformations. To this end, it features a geometric deformation augmenter that 1) enhances the robustness of global correction by addressing local deformations caused by non-rigid motion or geometric distortions, and 2) generates augmented data to improve the overall training process. UniMo is a hybrid model that uses both image intensities and shapes to achieve robust performance amid image appearance variations, and therefore generalizes effectively to multiple imaging modalities without requiring network retraining. We trained and tested UniMo to track motion in fetal magnetic resonance imaging, a particularly challenging application due to 1) the presence of both large rigid and non-rigid motion, and 2) wide variations in image appearance. We then evaluated the trained model, without retraining, on MedMNIST, lung CT, and BraTS datasets. The results show that UniMo surpassed existing motion correction methods in terms of accuracy, and notably enabled one-time training on a single modality while maintaining high stability and adaptability for inference across multiple unseen datasets. By offering a unified solution to motion correction, UniMo represents a significant advancement in medical imaging, especially in challenging applications with combined bulk motion and local deformations. The code is available at: https://github.com/IntelligentImaging/UNIMO
@@ I Introduction
Motion remains to be a challenge in medical imaging. For reliable imaging and image analysis, motion should be avoided, corrected, or compensated for, for reliable image analysis [1, 2, 3, 4]. Although classical image registration methods can be effective for motion correction, they often necessitate full execution of an optimization process on non-convex cost functions that only act as a surrogate measure of image alignment [5, 6, 7, 8, 9]. Thus, these techniques are typically limited by a narrow capture range, especially when addressing large motions. This constraint reduces their precision and applicability in challenging applications such as fetal MRI where subjects may move in wide angles [10, 11, 12]. To address these limitations, recent research has increasingly turned to deep learning based models for motion estimation/correction [13].
Deep learning-based motion correction techniques have seen widespread applications across various imaging modalities, with significant developments in two main areas: i) A major focus on Magnetic Resonance Imaging (MRI), including functional MRI (fMRI), T1-, T2-, and diffusion-weighted MRI, and quantitative MRI. These models are extensively studied due to their broad applicability and substantial impact in the field, covering a wide range of clinical scenarios. ii) Motion correction for other modalities, such as ultrasound, positron emission tomography (PET), and Computed Tomography (CT) where advancements and methodologies are also being explored.
Learning-based motion correction in MRI focuses on two key approaches: i) image-based motion correction and ii) techniques using k-space data. Image-based methods target motion correction by processing reconstructed MRI images with deep neural networks, such as Convolutional Neural Networks (CNNs) [11, 14] and other deep learning architectures [15, 16, 17, 18]. These approaches often employ convolutional encoder-decoder networks, leveraging downsampling and upsampling for feature extraction and reconstruction [19, 20]. Advanced models like U-Nets, recurrent neural networks, and transformers have been proposed to handle temporal dependencies and spatial relationships [21, 22, 18, 23, 15]. Additionally, prior-assisted methods improve correction by incorporating supplementary data such as multi-contrast or dynamic images, enhancing performance through architectures that integrate this information [24, 25]. These methods offer robust solutions by combining motion correction with other learning-based tasks, ultimately improving MRI reconstruction quality, e.g. [23].
@@
Despite significant advancements, state-of-the-art methods remain heavily constrained by the specific imaging datasets and modalities they are trained on and the types of motion they can handle, which limit their effectiveness in correcting motion in unseen imaging domains. While one approach [17] shows promise in handling motion for unseen modalities, it struggles with large motions in image pairs that are heavily contaminated by noise [39] and those that exhibit contrast changes. For instance, substantial intensity and contrast variations in images pose significant challenges, as current algorithms largely rely on image intensities while overlooking shape information. This limits generalization capacity of the trained models.
On the other hand, continuous advances in deep learning have enabled reliable and accurate segmentation and landmark detection, supporting their effective use in image registration. Compared to traditional intensity-based methods, shape- and keypoint-based approaches [14, 40, 41, 42] now provide robust alternatives, particularly in challenging imaging scenarios. Building on these developments, our joint optimization framework integrates shape cues to improve registration performance. This reflects a broader shift in the field toward leveraging geometric representations derived from learned features, which have demonstrated strong generalization across datasets and modalities [40]. Motivated by the need for generalizable and accurate motion correction across diverse imaging conditions, we introduce UniMo—a unified method that combines shape and intensity information to correct both bulk rigid motion and local deformations. UniMo capitalizes on the strengths of deep learning while avoiding the need for extensive retraining. Therefore, it offers a robust and efficient solution that is adaptable to various imaging conditions. By not imposing restrictive assumptions about motion types or ranges, UniMo ensures stable and accurate correction for a range of applications. Moreover, eliminating the retraining requirement reduces computational overhead and enables seamless integration into imaging workflows, which advances the state of the art in motion estimation and correction.
The main contributions of this work (UniMo) are:
  • UniMo is the first method for generalizable correction of both rigid and non-rigid motion without a need for retraining in a new domain.
  • UniMo offers motion correction of a pair of images with stable convergence in training and real-time inference.
@@ II-A Rigid Motion Estimation via Equivariant Filters
Rather than estimating the transformation function $Q$ directly in the original image space, efficient approaches have demonstrated that it can be computed using low-dimensional representations, such as key points or equivariant features of images [17]. In specific, equivariant filters have shown proven stability and accuracy in rigid image registration. The closed-form solution for both translation and rotation parameters is:
@@
where $\bar{S}$, $\bar{T}$ represent the low-dimensional representations of the source and target images respectively. $\Sigma V^{*}=\bar{S}\cdot\bar{T}^{T}, UU$ and $V^{*}$ are real orthogonal matrices, $\Sigma$ is a diagonal matrix with non-negative real numbers on the diagonal. The condition on the determinant of $\mathcal{R}$ in Eq. (2) is set to ensure it accurately reflects a rigid rotation.
@@ II-B Deformation Correction via Deformable Registration
where the distance function ${\rm Dist}(\cdot,\cdot)$ measures the dissimilarity between images. Commonly used distance functions include sum-of-squared difference of image intensities [43], normalized cross correlation [44], and mutual information [45, 41]. The regularization term $(\tilde{\mathcal{L}}\tilde{v}_{0},\tilde{v}_{0})$ enforces spatial smoothness of transformations with $\tilde{\mathcal{L}}$ being a symmetric and positive-definite differential operator in the Fourier domain. It converts a vector field $\tilde{v}$ to a momentum vector by $\tilde{m}=\tilde{\mathcal{L}}\tilde{v}$. The deformation $\psi_{1}$ corresponds to $\tilde{\psi}_{1}$ in Fourier space via the Fourier transform $\mathcal{F}(\psi_{1})=\tilde{\psi}_{1}$, or its inverse $\psi_{1}=\mathcal{F}^{-1}(\tilde{\psi}_{1})$.
@@
where $\star$ is a circular matrix-vector field auto-correlation [^1] Auto-correlation operates on zero-padded signals followed by truncating to the bandlimits in each dimension to ensure the output remains bandlimited.. $\tilde{\mathcal{K}}$ is a smoothing operator and $\tilde{\mathcal{D}}\tilde{v}$ is a tensor product with $\tilde{\mathcal{D}}(\mathbf{\xi})=i\sin(2\pi\mathbf{\xi})$ representing the Fourier frequencies of a central difference Jacobian matrix $D$. The operator $\tilde{\nabla}\cdot$ is the discrete divergence operator that is computed as the sum of the Fourier coefficients of the central difference operator $\tilde{D}$ along each dimension, i.e., $\tilde{\nabla}\cdot\mathbf{\xi}=\sum\limits_{j=1}^{d}i\sin(2\pi\mathbf{\xi}_{j})$. Since $\tilde{\mathcal{K}}$ on the left side of Eq. (4) is a low-pass filter that suppresses high frequencies in the Fourier domain, all operations are easy to implement in a truncated low-dimensional space by eliminating high frequencies.
@@ II-C Joint Optimization
$=\underbrace{\text{Dist}[S\circ Q(\mathcal{T},\mathcal{R}),T]+\text{Reg}(\mathcal{T},\mathcal{R})}_{\text{Rigid Motion Correction}}$
$+\underbrace{\text{Dist}(\mathbf{\hat{S}}\circ\psi_{1}(\tilde{v}_{0})),T)+(\tilde{\mathcal{L}}\tilde{v}_{0},\tilde{v}_{0})}_{\text{Deformation Correction}},$
@@
Such joint approaches have demonstrated enhanced accuracy and robustness for single image modalities [16, 15]. However, these frameworks fall short when image pairs exhibit significant intensity changes. The issue arises because both dissimilarity terms in Eq.(II-C) are highly dependent on the images themselves, increasing the likelihood that the optimization will be biased by intensity variations. This limitation motivated the development of a hybrid motion correction framework that integrates knowledge from both image intensities and shape to effectively reduce errors caused by image intensity variations.
@@ III Methodology
III-A Hybrid Motion Correction
[…]
@@ III-A Hybrid Motion Correction
After obtaining the hybrid rigid transformation $\boldsymbol{Q}$, for deformation estimation we estimate $\tilde{v}^{I}_{0}$ and $\tilde{v}^{G}_{0}$ (representing the transformation fields) in the spatial domain for images and segmentations respectively. The derivation of a hybrid velocity field is not the primary objective of this work. Instead, we formalize the hybrid objective function for optimization as follows,
@@
III-B Hybrid Rigid Transformation for Motion Correction
@@
[…]
@@
    The final hybrid rigid transformation matrix is
@@ III-C Network Design and Training
We develop a deep learning framework to estimate the objective expressed in Eq. (8). Our framework comprises two major sub-modules: i) A hybrid rigid motion correction neural network, parameterized by equivariant filters, to produce $\boldsymbol{Q}$; and ii) A hybrid deformation correction network, implemented using UNet, with deformable shape augmentation to estimate $\tilde{v}^{I}_{0}$ and $\tilde{v}^{G}_{0}$. Our framework is illustrated in Fig. 1. In the following sections, we provide a detailed description of the network architecture and the formulation of the network loss.
Figure 1: An illustration of the network architecture of our proposed motion correction learning framework, UniMO. Top: Motion correction is based on equivariant neural networks that take both images and segmentations as input. The low-dimensional spatial means of the images and segmentations are estimated simultaneously. An enhanced hybrid rigid transformation $Q$ is computed from both domains. The rigid loss function is calculated between the aligned and target images/labels. Bottom: A geometric shape augmenter is incorporated into the U-Net based neural networks. It takes both aligned and target images/segmentations and estimates the transformation fields for both domains using separate loss functions for segmentations and images.
Hybrid Rigid Motion Correction Equivariant Network
Let $\Theta=(\mathcal{T}_{I},\mathcal{R}_{I},\mathcal{T}_{G},\mathcal{R}_{G})$ represent the encoder parameters that learn rigid parameters from image and shape spaces, with $Q_{I}(\Theta)$ and $Q_{G}(\Theta)$ denoting the transformation functions yielded by the learned rigid parameters from the low-dimensional representations. Let $\boldsymbol{Q}$ denote the hybrid rigid transformation. The rigid correction loss is computed between the aligned outputs and targets. Note that CNNs or recurrent neural networks can also be used for extracting low-dimensional features from images [14, 11, 21]. However, such representations are limited in capturing the true nature of rigid transformations. In our approach, we employ an efficient method to compute rigid transformations within the equivariant neural network by calculating the spatial means of images [17]. We express the general formulation for the rigid motion network, which takes source images (or segmentations) and target images (or segmentations), as follows:
@@
In addition to estimating deformations using LDDMM [48], we provide a deep learning model that learns stationary velocity fields [54] while maintaining comparable model accuracy. Advanced predictive image registration models, including TransMorph [55] and DiffuseMorph [56], can be integrated into our framework. Analogous to our previous work [15], we extend UniMo to a spatio-temporal approach that incorporates time information to handle motion among data sequences.
@@
The objective function for the hybrid model is defined as: